OpenStudy (anonymous):

At Trent's garage, 3 cases of brake fluid last 4 weeks. Trent has 5 cases. How much more brake fluid should Trent order so that he has enough for 13 weeks? Round your answer up to the nearest whole number

OpenStudy (anonymous):

3/4 = x/13 -- 3 cases to 4 weeks = x cases to 13 weeks cross multiply (4)(x) = (13)(3) 4x = 39 x = 39/4 x = 9 3/4 cases 13 weeks would require 9 3/4 cases of break fluid. Trent has 5 cases. so.... 9 3/4 - 5 = 39/4 - 5 = 39/4 - 20/4 = 19/4 = 4 3/4 Trent would need 4 3/4 cases (along with his 5 cases) to have enough break fluid for 13 weeks.